Grade 10Algebra

if the first term of an AP is 2. the sum of fiveth term is equal to 1/4 of the next five term. find the sum of first 30 term?

Rohit Maurya
9 Years agoGrade 10

2 Answers

sukesh
9 Years ago
Let a be the first term. Then a =2.
 
sum of first 5 terms S5 = 5/2[2+2+(5-1)d]
                                = 10+10d
 
 
Sum of first 10 terms S10 = 10/2[2+2+(10-1)d]
                                     =  20+45d
 
For calculation of sum of next five terms is S10-S5 = 10+35d
 
 
sum of first five terms = ¼ sum of next 5
 
10+10d = 1/4[ 10+ 35d]
40+40d=10+35d
5d=-30
on simplification d=-6
 
no sum of first thirty terms = 30/2[2+2+(30-1)(-6)]
                                       =15[4-(29x6)]
                                        =15[-170]
                                        = -2550
Kushagra Madhukar
5 Years ago
Dear student,
Please find the answer to your question.
 
Let a be the first term. Then a =2.
sum of first 5 terms S5 = 5/2[2+2+(5-1)d]
                                = 10+10d
Sum of first 10 terms S10 = 10/2[2+2+(10-1)d]
                                     =  20+45d
For calculation of sum of next five terms is S10-S5 = 10+35d
sum of first five terms = ¼ sum of next 5
10+10d = 1/4[10+ 35d]
40+40d=10+35d
5d = – 30
on simplification d = – 6
no sum of first thirty terms = 30/2[2+2+(30–1)(–6)]
                                       =15[4 – (29x6)]
                                        =15[ – 170]
                                        = – 2550
 
Thanks and regards,
Kushagra